OpenStudy (micheal1234):

Adele’s lawn has an area of (x² – 5x – 6) square feet and a length of (x + 1) feet. What is the width of the lawn in feet?

ganeshie8 (ganeshie8):

Is his lawn a rectangle ?

OpenStudy (micheal1234):

dont know

ganeshie8 (ganeshie8):

Let's assume it is a rectangle

ganeshie8 (ganeshie8):

Remember the formula for area of rectangle?

OpenStudy (micheal1234):

sorry it took so long to get back to you computer went on the frights do you think you can set it up for me

OpenStudy (micheal1234):

ganeshie8

ganeshie8 (ganeshie8):

I'm trying to help you set it up

ganeshie8 (ganeshie8):

Again, what's the formula for area of rectangle ?

OpenStudy (micheal1234):

it is A=wl

ganeshie8 (ganeshie8):

Good. Substitute the given expressions in that formula

OpenStudy (micheal1234):

and what is that

ganeshie8 (ganeshie8):

Look at the question You are given expressions for area and length

ganeshie8 (ganeshie8):

A = w*l

OpenStudy (micheal1234):

so would i divide the to

ganeshie8 (ganeshie8):

Plugging the given expressions in above formula : \[x^2-5x-6 = w*(x+1)\]

ganeshie8 (ganeshie8):

Yes. Divide (x+1) both sides

ganeshie8 (ganeshie8):

Do you know how to factor the left side quadratic?

OpenStudy (micheal1234):

so (x-6)

OpenStudy (micheal1234):

you still there

ganeshie8 (ganeshie8):

Excellent!